### v1

@@ 5 Conclusion
Our work introduces the Unified-EGformer, addressing mixed exposure challenges in images with a novel transformer model. Through specialized local and global refinement alongside guided attention, it demonstrates superior performance across various scenarios. Its lightweight architecture makes it suitable for real-time applications, advancing the field of image enhancement and restoration. Enhancing Unified-EGFormer could involve refining the attention mechanism to become color independent to diminish the influence of color artifacts in the enhanced output. Additionally, exploring the integration of lightweight state space models [14], with bi-exposure guidance offers promising avenues for further optimizing the network for efficiency and performance in image enhancement tasks.
@@
6 Future Work & Discussion
The challenges presented by mixed exposure in images are not just specific to any single methodology but are a broader issue within the field of image enhancement. As illustrated in [Fig. 4], both overexposure and underexposure present unique challenges. Overexposure often results in the loss of critical details such as edges and contours, while underexposure can hide essential information in darkness. These scenarios make effective image reconstruction a complex task.
A further complication arises when considering the standard for evaluation—what exactly constitutes an ideal ground truth in the context of mixed exposure? Often, ground truth images themselves may lack details in over-exposed areas, complicating the assessment of enhancement algorithms as can be seen in [Fig. 4]. This highlights a significant gap in our current understanding and capabilities, emphasizing the need for advancements that can precisely discern and correct varying degrees of exposure while preserving the integrity of the image details.
The field of mixed exposed is not new, however has limited exploration and yet far away to successfully solve the problem. Potential solutions such as Integrating Kolmogorov-Arnold Networks (KANs) [40] in place of MLP blocks in transformers can yield more efficient and explainable models. KANs’ adaptive activation functions can better distinguish and process overexposed and underexposed regions, enhancing image quality.
While current methods address many aspects of image quality, there is still room for improvement in creating color-independent ground truths for guided attention map generators. Color-independent techniques prevent issues such as greens appearing yellow due to color influence, ensuring accurate hue representation and processing. Understanding and correcting these color dynamics involve using attention map generators to identify inconsistencies.
@@ 7 Model Design Components
This extended material provides with more details in our model design components.
@@ 7.1 Algorithm
The details presented in the training are shown in the algorithm below:
@@
Manual pixel-level Otsu thresholding mask:
@@ 6 $\mathbf{D}_{i}\leftarrow$ Operation($\mathbf{D}$)
EGFormer excutes Training$(\mathbf{D},\mathbf{M}_{g},\mathbf{I}_{g},\mathbf{W}_{LEB},\mathbf{W}_{GEB},\mathbf{W}_{EAF},E)$:
7 for each epoch $e=1,2,...,E$ do
@@ 7.2 A-MSA
[Eq. 7.3] The DGFN mechanism involves two parallel pathways that process the input features $m\in\mathbf{M}_{g}$ through distinct gating mechanisms, utilizing GELU activations ($\phi$) and element-wise products. Each path applies a sequence of convolutional transformations to enrich local context, comprising a $1\times 1$ convolution followed by a $3\times 3$ depth-wise convolution. Moreover, depth-wise convolutions facilitate spatial information processing while significantly reducing computational costs, making them particularly suitable for developing a light-weight environment without compromising performance. Following Wang et al.’s, [57] approach, the outputs of these parallel pathways are then merged using element-wise summation, allowing for a comprehensive feature refinement that incorporates both detailed and global information.
@@ 7.4 Model Training
For LOL-v1 [60], we adopt the standard split provided, comprising images of dimensions $400\times 600$, with 500 images allocated for both training and evaluation. Similarly, for LOL-v2 [60], we utilize the real-world dataset for both training and evaluation, consisting of images of size $400\times 600$, with 689 and 100 images in the respective splits. We adopt patch sizes of $256\times 256$ and $324\times 324$ for training across both LOL-v1 and LOL-v2 datasets. Additionally, we incorporate the MIT-Adobe FiveK [3] (FiveK) dataset for comparative analysis with other baseline methods. The FiveK dataset comprises 5,000 image pairs, each meticulously retouched by experts (A - E), with the best-performing result reported. To align with established practices [11, 2], we crop the training images to $256\times 256$ dimensions, subsequently fine-tuning over $324\times 324$ and $512\times 512$ patches, while the test images are resized to a maximum dimension of 512 pixels. Additionally, we incorporate the MSEC [2] dataset, which features realistic underexposure and overexposure scenarios derived from the MIT-FiveK dataset. Our experimental setup involves 17,675 training images, 750 for validation, and 5,905 for testing, adhering to the general experimental protocol of our baselines. Furthermore, we utilize the SICE [4] dataset, employing the SICE-v2 variant, a resized version of the original dataset, following the approach outlined in [21, 20]. Lastly, to comprehensively evaluate mixed-exposure scenarios, we assess our model’s performance on two variants of the SICE dataset: SICE-Grad and SICE-Mix [74], which feature mixed under and overexposure regions within single images.
To facilitate network training, we employ the cosine annealing schedule along with the Cosine-Annealing learning rate scheduler, incorporating 15 warm-up epochs and setting eta_min to $51e-$ 5. Additionally, we utilize the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, along with an epsilon value of $81e-$ 8. Our initial learning rate is set to $41e-$ 4 for pretraining purposes. Furthermore, we adopt a batch size of 4 and a weight decay of $41e-$ 4 to stabilize training and improve model performance.
@@ 8.1 Generalization in the Wild
Our model demonstrates impressive visual qualitative results on the BAID dataset. Initially trained on the LOLv2 dataset, the model was finetuned using only 300 randomly selected images from the BAID train dataset, out of the 3000 available backlit images. These BAID images were originally downsampled images and were directly used from Liang et al., [36]. The backlit images in BAID present a challenging scenario due to uneven exposure, typically featuring underexposure in the foreground and overexposure in the background. Traditional approaches often struggle with these conditions, but Unified-EGformer, particularly with our Global Enhancement Block (GEB) combined with the Exposure Aware Fusion Block (EAF), excel by exploiting contextual features. Despite the limited finetuning data, our model significantly enhances backlit images, effectively generalizing to both indoor and outdoor scenes under diverse backlit conditions.
Figure 10: U-EGformer’s output from ‘Local Echancemnet Block (LEB)’ and ‘Global Enhancement Block (GEB) + Exposure Aware Fusion Block (EAF)’ enhancement modules compared with Ground-truth image on BAID set.
@@
Figure 13: Showing the Global+EAF block’s significance when adverse lighting brings artifacts in the reconstructed image. First column: ‘Local Echancemnet Block (LEB)’ and second column: ‘Global Enhancement Block (GEB) + Exposure Aware Fusion Block (EAF)’ enhancement modules compared with third column: Ground-truth image on BAID test set.

### v2

@@ 5 Conclusion
Our work introduces Unified-EGformer, addressing mixed exposure challenges in images with a novel transformer model. Through specialized local and global refinement alongside guided attention, it demonstrates superior performance across various scenarios. Its lightweight architecture makes it suitable for real-time applications, advancing the field of image enhancement and restoration. Unified-EGformer could be enhanced further by refining the attention mechanism to become color independent to diminish the influence of color artifacts in the enhanced output. Additionally, exploring the integration of lightweight state space models [gu2024mamba, adhikarla2024expomamba], with bi-exposure guidance offers promising avenues for further optimizing the network for efficiency and performance in image enhancement tasks.
@@
6 Discussion & Future Work
The challenges presented by mixed exposure in images are not just specific to any single methodology but are a broader issue within the field of image enhancement. As illustrated in Fig. 4, both overexposure and underexposure present unique challenges. Overexposure often results in the loss of critical details such as edges and contours, while underexposure can hide essential information in darkness. These scenarios make effective image reconstruction a complex task.
A further complication arises when considering the standard for evaluation—what exactly constitutes an ideal ground truth in the context of mixed exposure? Often, ground truth images themselves may lack details in over-exposed areas, complicating the assessment of enhancement algorithms as can be seen in Fig. 4. This highlights a significant gap in our current understanding and capabilities, emphasizing the need for advancements that can precisely discern and correct varying degrees of exposure while preserving the integrity of the image details.
The field of mixed exposure is not new; however it has seen limited exploration and is thus far from successfully solving the problem. Future research could consider integrating Kolmogorov-Arnold Networks (KANs) [liu2024kan] in place of MLP blocks in transformers as they can yield more efficient and explainable models. KANs’ adaptive activation functions can better distinguish and process overexposed and underexposed regions, enhancing image quality.
While current methods address many aspects of image quality, there is still room for improvement in creating color-independent ground truths for guided attention map generators. Color-independent techniques prevent issues such as greens appearing yellow due to color influence, ensuring accurate hue representation and processing. Understanding and correcting these color dynamics may require using attention map generators to identify inconsistencies.
@@ 7 Model Design Components
This extended material provides more details of our model design components.
@@ 7.1 Algorithm
The details presented in the training are shown in Algorithm 1 below:
@@
1
1exManual pixel-level Otsu thresholding mask:
@@ 8 $\mathbf{D}_{i}\leftarrow$ Operation($\mathbf{D}$)
9
[…]
10 for each epoch e=1,2,,Ee=1,2,...,E do
@@ 7.2 A-MSA
The DGFN mechanism involves two parallel pathways that process the input features $m\in\mathbf{M}_{g}$ through distinct gating mechanisms, utilizing GELU activations ($\phi$) and element-wise products. Each path applies a sequence of convolutional transformations to enrich local context, comprising a $1\times 1$ convolution followed by a $3\times 3$ depth-wise convolution. Moreover, depth-wise convolutions facilitate spatial information processing while significantly reducing computational costs, making them particularly suitable for developing a light-weight environment without compromising performance. Following the approach in Wang et al. [wang2023ultra], the outputs of these parallel pathways are then merged using element-wise summation, allowing for a comprehensive feature refinement that incorporates both detailed and global information.
@@ 7.4 Model Training
For LOL-v1 [WeiWY018], we adopt the standard split provided, comprising images of dimensions $400\times 600$, with 500 images allocated for both training and evaluation. Similarly, for LOL-v2 [WeiWY018], we utilize the real-world dataset for both training and evaluation, consisting of images of size $400\times 600$, with 689 and 100 images in the respective splits. We adopt patch sizes of $256\times 256$ and $324\times 324$ for training across both LOL-v1 and LOL-v2 datasets. Additionally, we incorporate the MIT-Adobe FiveK [fivek] dataset for comparative analysis with other baseline methods. The FiveK dataset comprises 5,000 image pairs, each meticulously retouched by experts (A - E), with the best-performing result reported. To align with established practices [Cui_2022_BMVC, afifi2021learning], we crop the training images to $256\times 256$ dimensions, subsequently fine-tuning over $324\times 324$ and $512\times 512$ patches, while the test images are resized to a maximum dimension of 512 pixels. Additionally, we incorporate the MSEC [afifi2021learning] dataset, which features realistic underexposure and overexposure scenarios derived from the MIT-FiveK dataset. Our experimental setup involves 17,675 training images, 750 for validation, and 5,905 for testing, adhering to the general experimental protocol of our baselines. Furthermore, we utilize the SICE [Cai2018deep] dataset, employing the SICE-v2 variant, a resized version of the original dataset, following the approach outlined in [huang2022exposure-consistency, huang2023learning]. Lastly, to comprehensively evaluate mixed-exposure scenarios, we assess our model’s performance on two variants of the SICE dataset: SICE-Grad and SICE-Mix [zheng2022low], which feature mixed under and overexposure regions within single images.
To facilitate network training, we employ the cosine annealing schedule along with the Cosine-Annealing learning rate scheduler, incorporating 15 warm-up epochs and setting eta_min to 1e51e-5. Additionally, we utilize the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, along with an epsilon value of 1e81e-8. Our initial learning rate is set to 1e41e-4 for pretraining purposes. Furthermore, we adopt a batch size of 4 and a weight decay of 1e41e-4 to stabilize training and improve model performance.
@@ 8.1 Generalization in the Wild
Our model demonstrates impressive visual qualitative results on the BAID dataset. Initially trained on the LOLv2 dataset, the model was finetuned using only 300 randomly selected images from the BAID train dataset, out of the 3000 available backlit images. These BAID images were originally downsampled images and were directly used from Liang et al. [liang2023iterative]. The backlit images in BAID present a challenging scenario due to uneven exposure, typically featuring underexposure in the foreground and overexposure in the background. Traditional approaches often struggle with these conditions, but Unified-EGformer, particularly with our Global Enhancement Block (GEB) combined with the Exposure Aware Fusion Block (EAF), excel by exploiting contextual features. Despite the limited finetuning data, our model significantly enhances backlit images, effectively generalizing to both indoor and outdoor scenes under diverse backlit conditions.
Figure 9: U-EGformer’s output from ‘Local Enhancement Block (LEB)’ and ‘Global Enhancement Block (GEB) + Exposure Aware Fusion Block (EAF)’ enhancement modules compared with Ground-truth image on BAID set.
@@
Figure 12: Showing the Global+EAF block’s significance when adverse lighting brings artifacts in the reconstructed image. First column: ‘Local Enhancement Block (LEB)’ and second column: ‘Global Enhancement Block (GEB) + Exposure Aware Fusion Block (EAF)’ enhancement modules compared with third column: Ground-truth image on BAID test set.